# Characterizing the Complexity of Social Robot Navigation Scenarios

### v1

@@ Abstract
Social robot navigation algorithms are often demonstrated in overly simplified scenarios, prohibiting the extraction of practical insights about their relevance to real-world domains. Our key insight is that an understanding of the inherent complexity of a social robot navigation scenario could help characterize the limitations of existing navigation algorithms and provide actionable directions for improvement. Through an exploration of recent literature, we identify a series of factors contributing to the complexity of a scenario, disambiguating between contextual and robot-related ones. We then conduct a simulation study investigating how manipulations of contextual factors impact the performance of a variety of navigation algorithms. We find that dense and narrow environments correlate most strongly with performance drops, while the heterogeneity of agent policies and directionality of interactions have a less pronounced effect. This motivates a shift towards developing and testing algorithms under higher-complexity settings.
@@ I Introduction
Although it is a time of rich research activity in the social navigation field [36, 15, 54], researchers often evaluate their methods with strong assumptions that ultimately do not transfer to the real world. Recent surveys on social robot navigation [36, 15, 54] have highlighted systematic problems that stem from the lack of concrete benchmarking protocols. Much of the literature considers sparse or slowly navigating human crowds, which give rise to relatively non-interactive encounters that are easy to handle by classical navigation algorithms. Hence, it remains unclear whether modern approaches are truly effective. However, it is not straightforward to define a rigorous benchmark because there is a lack of community consensus over appropriate experiment design, which would need to capture a multitude of aspects such as the spatial arrangement and roles of all agents, the briefing that human users receive before the experiment, robot design specifications, metrics evaluating robot performance and human impressions, etc.
@@
In this paper we argue that to establish effective benchmarks for social robot navigation, the research community should better understand and control the dimensions of problem complexity. Fields like theoretical computer science, learning theory, and motion planning seek to establish clear definitions of complexity [2, 65, 13] that help define the limits of existing algorithms and motivate directions for research. However, robotics is a synthetic science [25], combining many elements from other fields, which makes it harder to formulate similarly succinct and relevant definitions for complexity. Social robot navigation [36] is no exception to this complication, as it requires the consideration of multiple disciplines, including motion planning, machine learning, control, and human factors, among others. We choose to define complexity using an empirical approach: a dimension of Complexity is a factor that causes a drop in Performance of an automated method relative to average humans (Fig. 1). Here, Performance refers to any relevant metric accepted by the community.
Figure 1: We investigate how different parameters defining a social robot navigation scenario impact the performance of recent algorithms. We show that while many algorithms successfully handle a wide range of scenarios, their performance drops at a region of Complexity that is representative of real-world situations that humans effectively handle [19].
@@
To instantiate this analytical framework, we review the evaluation practices of recent literature and extract a set of factors contributing to the complexity of a social robot navigation scenario, distinguishing contextual factors from robot-related ones. We then conduct an extensive simulated study to understand how robot-independent, contextual factors impact algorithm performance. We find that Density and Environment Geometry have the strongest correlation with performance decline across a wide range of algorithms. We also find that the scenarios most frequently explored in prior work tend to be lower complexity, and allow simple, reactive methods to perform effectively. Our findings underscore the need for algorithms that explicitly account for diverse, high Complexity scenarios, and benchmarks that explore performance in scenarios that have high Complexity across multiple factors.
@@ II Related Work
Recent surveys on social robot navigation [36, 15, 54, 16] emphasize the need for a standardization of benchmarking practices. Mavrogiannis et al. [36] observe the lack of rigorous statistical characterization of robot performance and human impressions, and emphasize the need for better repeatability. Francis et al. [15] collect a list of principles that a social navigation system should take into account, and propose guidelines for evaluating them. Gao and Huang [16] delve into the characterization of methodologies adopted by researchers and practitioners. Singamaneni et al. [54] provide a taxonomy of the literature, capturing aspects such as the robot type, the algorithmic approaches used, and the social navigation context. This work complements prior surveys by analyzing the hardness of the scenarios employed in the evaluation practices of published approaches. By understanding the explored territory of complexity, we aspire to provide directions for integrating more complex settings in benchmarking practices that more closely resemble what a robot would experience in the real world.
@@
A major challenge in standardizing benchmarking in social navigation lies in the design of experiments, and specifically in balancing repeatability and emergence of natural interactions. Mavrogiannis et al. [36] classify real-world experiments into three classes: demonstrations (e.g., [57, 28, 29]), lab studies (e.g., [37, 35, 46, 42, 31]), and field studies (e.g., [59, 58, 6]), advocating the need for the extraction of statistical insights across social navigation experiments. However, the cost of conducting real-world experiments has motivated the experimentation within virtual environments [5]. Borrowing tools from graphics and the simulation of crowds [64, 20], several works have developed photorealistic simulators [21, 5, 62], while others have focused on finding quantitative metrics that capture aspects of performance previously only understood through qualitative metrics [66, 34, 53]. However, as demonstrated by Fraichard and Levesy [14], simulation is often introduced in evaluation practices with shortcomings. Common assumptions in simulated evaluations include the robot being invisible to humans, humans being simulated as homogeneous overly submissive agents, the scenarios considered being generated at random, and often in open spaces with no walls or obstacles. Such assumptions result in surprising observations, such as that a “blind” robot moving straight can outperform state-of-the-art agents [36]. Motivated by these observations, we contribute an experiment design that gives rise to more realistic crowd conditions by e.g., designing heterogeneous agents, and considering the manipulation of contextual variables like the environment geometry, the crowd directionality and density.
@@
Another challenge in designing benchmarks is the lack of understanding of what makes a social navigation scenario hard. In prior work, the crowd density [59, 36, 55] has often been used as a proxy for complexity. However, density on its own does not capture the complexity of the motion coupling between closely interacting agents. To capture that, metrics like Path Irregularity [18] and the Topological Complexity Index [12, 34] quantify aspects of geometric and topological richness of agents’ trajectories. Our work is relevant to these approaches but it seeks to extract a more fundamental understanding of how the parameters defining a social navigation scenario impact its hardness. We are inspired by the Complexity definitions that more theoretical fields employ [2, 65, 13]. While our investigation is empirical, we see this as a first step towards the definition of a formal representation of Complexity.
@@ III Factors of Complexity in the Literature
To approach a characterization of the Complexity of a social navigation scenario, we investigated the benchmarking practices of prior work, searching for relevant papers using the keywords “social” and “crowd” in proceedings from the ICRA, IROS, CoRL, and RSS conferences from 2015 to 2024. Additionally, we reviewed work cited in recent surveys [36, 15, 54, 16]. We organized the characteristics of evaluation methodologies and found a helpful breakdown of the factors characterizing the evaluation of each scenario could be through two categories: contextual (see Fig. 2) and robot-related factors. We tabulate the results in Table I.
@@ III-A Contextual Factors
By contextual, we refer to factors are robot agnostic, i.e., they are related to the environment the robot is deployed in.
@@
Directionality. The directions in which humans encounter the robot contribute to the difficulty of the collision-avoidance task [67]. We identified four cases: Passing: Agents move parallel, paths do not intersect; Crossing: Agents move perpendicular, paths intersect; Random: Agent starts and goals are randomly sampled; Circle Crossing: Agent start and goals are sampled on opposite sides of the circumference of a circle.
@@
Policy Mixture. The way humans navigate next to the robot is another complicating factor. Cooperative agents assume partial responsibility for collision avoidance, which simplifies the robot’s task. In contrast, when rushing, being distracted, or changing intentions, humans may pose greater challenges to a robot [46]. We found that most real-world studies instruct participants to navigate cooperatively, and most simulation-based studies make use of cooperative crowd simulators like Social Force [20] and ORCA [63].
@@ III-B Robot Factors
Maximum Speed. Consistently, the maximum speed of the robot is much lower than a natural human walking speed [46], due to hardware or safety constraints. When reviewing supplementary video material, however, we observed that this often resulted in the humans moving to avoid the robot, while the robot maintained a nearly straight trajectory.
@@
Viewpoint. We differentiate between overhead, in which sensors have a bird’s eye view of the scene, mitigating issues such as occlusions, and egocentric, in which sensors are placed on the robot.
@@ IV Investigating the Complexity of Social Robot Navigation Scenarios
We propose a definition of a Social robot navigation scenario based on parameters related to the factors of Complexity identified in Sec. III. We then describe an experiment design that investigates the performance of a variety of navigation algorithms under different scenarios.
@@ IV-A Social Navigation Scenario
Consider a robot navigating next to $n\geq 1$ human agents in a workspace $\mathcal{W}\subseteq\mathbb{R}^{2}$ with a set of static obstacles $\mathcal{W}_{obs}\subseteq\mathcal{W}$. The robot starts from an initial configuration $s_{R}$ and moves towards a goal $g_{R}$ by following a policy $\pi_{R}$ whereas humans are navigating from their initial configurations $s_{i}$ towards their goals, $g_{i}$ by following a policy $\pi_{i}$, $i\in\mathcal{N}$; agents’ goals are unknown to one another. The robot occupies an area $A_{R}\in\mathcal{W}$, and each human occupies an area $A_{i}\in\mathcal{W}$. The objective of the robot is to reach its destination while avoiding collisions with static obstacles and abiding by social norms. We define a social navigation scenario as a tuple:
$\mathcal{S}=\left(n,A_{R},A_{i:n},s_{R},g_{R},s_{i:n},g_{i:n},\pi_{i:n},\mathcal{W}_{obs}\right)\mbox{.}$ (1)
@@ IV-B Experiment Design
Scenario configurations. We first define a base scenario $\mathcal{S}_{b}$ defining with $n=15$, $A_{R},A_{i:n}=\pi(0.3)^{2}$, $s_{R}=(5,1)$, $g_{R}=(5,9)$, $\pi_{i:n}=ORCA,SFM,v_{pref}=1.0m/s$, $\mathcal{W}_{obs}=\{[0,w]\times[0,l]\}^{\mathsf{c}}=\{[0,10]\times[0,10]\}^{\mathsf{c}}$; $s_{i:n},g_{i:n}$ are sampled from passing and crossing. We then modify a single variable in each experiment, from low to high intensity:
@@
  • Policy Mixture: {SFM only, ORCA only, Mix 1, Mix 2, Mix 3}, where Mix 1 contains 8 ORCA and 7 SFM agents, Mix 2 contains 5 ORCA, 5 SFM, 2 CV, 3 Static agents, and Mix 3 contains 4 ORCA, 4 SFM, 4 CV, 3 Static agents, respectively. We add increasing numbers of inattentive agents at higher intensity to model Complex real-world scenarios in which the robot must navigate among both cooperative and uncooperative agents simultaneously [33].
  • Environment Width: {4.5, 4.0, 3.5, 3.0, 2.5, 2.0, 1.5}. , which reflects the reported width of most hallways (e.g., [24, 71, 42]) in Table. I.
@@
Refer to caption   Refer to caption   Refer to caption
a Density
Refer to caption   Refer to caption   Refer to caption
b Width
Refer to caption   Refer to caption   Refer to caption
c Directionality
Refer to caption   Refer to caption   Refer to caption
d Agent Policy Mixture
Figure 2: Left to right: Scenarios of increasing complexity for each of the complexity factors considered (d-p). In the leftmost Directionality figure green agents are crossing and blue agents are passing, while different colors represent different policies in Policy Mixture.
@@
  • Social GAN Model Predictive Control (MPC-SGAN) [17]: An approach integrating a recent crowd motion prediction model into an MPC framework, using the approach and implementation of Poddar et al. [46].
@@
Metrics. Based on literature  [16, 36, 15], we use the following metrics to measure performance, collected only from Successful trials:
@@
Hypotheses. We expect that scenarios of higher Complexity will pose greater navigation challenges, and this will be reflected in significant performance drops across all algorithms. Additionally, as Density is often used as a proxy for Complexity as a whole, we anticipate that it will have the strongest correlation with performance drop. Based on these expectations, We formulate the following hypotheses:
  • H1. Increasing the intensity of each of the four complexity factors (Density, Directionality, Environment Geometry, Policy Mixture) independently decreases performance with respect to (wrt) collected metrics.
@@ IV-C Implementation Details
We generate 500 scenarios for each condition within each experiment and report metrics as averages across all 500 trials. We fix the random seed to ensure each method experiences the same scenarios. To simulate continuous crowd motion, agents are randomly assigned new collision free starts and goals upon reaching their goal. For the MPC methods, we used the formulation of Poddar et al. [46]. For MPC-SGAN and MPPI-SGAN, we use a checkpoint pretrained on the Zara portion of the UCY dataset [26]. For the ORCA and SFM methods, we used the default parameter settings. The reactive planner uses the same action space of Chen et al. [8], and selects the collision free action that minimizes its distance to the goal in the following timestep. All methods handle static obstacle collisions by setting the component of the action in the direction of the colliding obstacle to 00. For specific details of the training procedure for RGL, and parameter settings and tuning for the MPC and MPPI methods, as well as our enhanced version of CrowdNav, see our GitHub site at https://github.com/fluentrobotics/ComplexityNav. Videos of our experiments can be found at https://youtu.be/-ir12VoSCkY.

### v2

@@ Abstract
Social robot navigation algorithms are often demonstrated in overly simplified scenarios, prohibiting the extraction of practical insights about their relevance to real-world domains. Our key insight is that an understanding of the inherent complexity of a social robot navigation scenario could help characterize the limitations of existing navigation algorithms and provide actionable directions for improvement. Through an exploration of recent literature, we identify a series of factors contributing to the complexity of a scenario, disambiguating between contextual and robot-related ones. We then conduct a simulation study investigating how manipulations of contextual factors impact the performance of a variety of navigation algorithms. We find that dense and narrow environments correlate most strongly with performance drops, while the heterogeneity of agent policies and directionality of interactions have a less pronounced effect. Our findings motivate a shift towards developing and testing algorithms under higher-complexity settings.
Index Terms:
Human-Aware Motion Planning, Human-Centered Robotics, Autonomous Vehicle Navigation.
@@ I Introduction
Recent surveys on social robot navigation [1, 2, 3] have highlighted that strong assumptions in typical evaluation practices prohibit transfer to real-world domains. Much of the literature considers sparse, slowly navigating human crowds, giving rise to non-interactive human-robot encounters that are easily handled by classical navigation algorithms. Hence, the benefits of many modern approaches are not transparent. This underscores the need for standardized benchmarks; however, defining benchmarks for this domain requires community consensus over appropriate experimental design. This would need to capture several aspects, such as the spatial arrangement and roles of agents, the briefing that human users receive before the experiment, robot design specifications, metrics evaluating robot performance and human impressions, etc.
@@
In this paper we argue that to establish effective benchmarks for social robot navigation, the research community should better understand and control the dimensions of problem Complexity. Clear definitions of Complexity in fields like theoretical computer science [4], learning theory [5], and motion planning [6] enable a characterization of limitations in existing algorithms and motivate future research directions. However, robotics is a synthetic science [7], combining many elements from other fields, which makes it harder to formulate similarly succinct and relevant definitions for Complexity. Social robot navigation [1] is no exception to this complication, as it lies at the intersection of multiple research areas, including motion planning, machine learning, control, and human modeling, among others. In this work, we approach Complexity through an empirical lens: a dimension of Complexity is a factor that causes a drop in Performance of an automated method (Fig. 1). Here, Performance is measured with respect to relevant criteria accepted by the community [1].
Figure 1: Humans can seamlessly handle a wide range of crowd navigation scenarios [8]. In contrast, social robot navigation algorithms struggle to handle scenarios with realistic levels of Complexity [9]. In this paper, we show that high-Complexity scenarios are underexplored in the literature. We demonstrate principal factors contributing to poor navigation performance, and argue for the definition of benchmarks that account for their impact.
@@
To instantiate our framework, we review the evaluation practices of recent literature and extract a set of factors contributing to the Complexity of a social robot navigation scenario, distinguishing between contextual and robot-related factors. We then conduct an extensive simulated study to understand how robot-independent, contextual factors impact algorithmic performance. We find that Density and Environment Geometry have the strongest correlation with performance decline across a wide range of algorithms. We also find that the scenarios most frequently explored in prior work tend to be of lower Complexity, and allow simple, reactive methods to perform effectively. Our findings underscore the need for algorithms that explicitly account for diverse, high-Complexity scenarios, and benchmarks that explore performance in scenarios that have high Complexity across multiple factors.
@@ II Related Work
Recent surveys on social robot navigation [1, 2, 3, 10] emphasize the need for a standardization of benchmarking practices. Mavrogiannis et al. [1] observe the lack of rigorous statistical characterization of robot performance and human impressions, and emphasize the need for better repeatability. Francis et al. [2] collect a list of principles that a social navigation system should take into account, and propose guidelines for evaluating them. Gao and Huang [10] delve into the evaluation methodologies adopted by researchers and practitioners. Singamaneni et al. [3] provide a taxonomy of the literature, capturing aspects such as the robot type, the algorithmic approaches used, and the social navigation context. This work complements prior surveys by analyzing the Complexity of the scenarios employed in the evaluation practices of published approaches. By understanding the explored territory of Complexity, we provide directions for integrating more Complex settings in benchmarking practices that more closely resemble what a robot would experience in the real world.
@@
A major challenge in standardizing benchmarking in social navigation lies in the design of experiments, and specifically in balancing repeatability and emergence of natural interactions. Mavrogiannis et al. [1] classify real-world experiments into three classes: demonstrations [11, 12, 13], lab studies [14, 15, 16, 17, 18], and field studies [9, 19, 20], advocating for the extraction of statistical insights across social navigation experiments. However, the cost of conducting real-world experiments has motivated experimentation within virtual environments [21]. Borrowing tools from graphics and crowd simulation [22, 23], several works have developed photorealistic simulators [24, 21, 25], while others have focused on finding quantitative metrics that capture aspects of performance previously only understood through qualitative metrics [26, 27, 28]. However, as demonstrated by Fraichard and Levesy [29], simulation is often introduced in evaluation practices with shortcomings. Common assumptions in simulated evaluations include the robot being invisible to humans, humans being simulated as homogeneous overly submissive agents, the scenarios considered being generated at random, and the scenarios containing no walls or obstacles. Such assumptions result in surprising observations, such as that a “blind” robot moving straight can outperform state-of-the-art agents [1]. Motivated by these observations, we contribute an experimental design that gives rise to more realistic crowd conditions by considering the manipulation of contextual variables including the Environment Geometry, the agent Policy Mixture, the crowd Directionality, and Density.
@@
Another challenge in designing benchmarks is the lack of understanding of what makes a social navigation scenario Complex. In prior work, the crowd Density [9, 1, 30] has often been used as a proxy for Complexity. However, Density on its own does not capture the Complexity of the motion coupling between closely interacting agents. To capture that, metrics like Path Irregularity [31] and the Topological Complexity Index [32, 27] quantify aspects of geometric and topological richness of agents’ interactions. Our work is relevant to these approaches, however it seeks to extract a more fundamental understanding of how the parameters defining a social navigation scenario impact its Complexity. We are inspired by the Complexity definitions of theoretical fields [4, 5, 6]. While our investigation is empirical, we see it as a first step towards defining a formal representation of Complexity.
@@ III Factors of Complexity in the Literature
We first identified factors of Complexity in the literature by investigating benchmarking practices in papers from the ICRA, IROS, CoRL, and RSS conferences from 2015 to 2024, using the keywords “social” and “crowd”. Additionally, we reviewed work cited in recent surveys [1, 2, 3, 10]. We studied a variety of evaluation methodologies and found a helpful breakdown of the factors characterizing the evaluation of each scenario could be through two categories: contextual (see Fig. 2) and robot-related factors. We tabulate the results in Table I.
@@ III-A Contextual Factors
By contextual, we refer to factors are robot-agnostic, i.e. they are related to the environment the robot is deployed in.
@@
Directionality. The directions in which humans encounter the robot contribute to the difficulty of the collision-avoidance task [62]. We identified four cases– Passing: Agents move parallel, paths do not intersect; Crossing: Agents move perpendicular, paths intersect; Random: Agent starts and goals are randomly sampled; Circle Crossing: Agent start and goals are sampled on opposite sides of the circumference of a circle.
@@
Policy Mixture. The behavior with which each agent co-navigates affects the evolution of a scenario. Cooperative agents assume partial responsibility for collision avoidance, which simplifies the robot’s task. In contrast, when rushing, being distracted, or changing intentions, humans may pose greater challenges to a robot [16]. We found that most real-world studies instruct participants to navigate cooperatively, and most simulation-based studies make use of cooperative crowd simulators like Social Force [23] and ORCA [63].
@@ III-B Robot Factors
Maximum Speed. Consistently, the maximum speed of the robot is much lower than a natural human walking speed [16], due to hardware or safety constraints. When reviewing supplementary video material, however, we observed that this often resulted in the humans moving to avoid the robot, while the robot itself performed little to no socially aware maneuvers.
@@
Viewpoint. We differentiate between overhead, in which sensors have a bird’s eye view, mitigating issues like occlusions, and egocentric, in which sensors are on the robot.
@@ IV Investigating the Complexity of Social Robot Navigation Scenarios
We propose a definition of a Social robot navigation scenario based on parameters that concisely capture the contextual factors in Sec. III. We then describe an experiment design that investigates the performance of a variety of navigation algorithms under different scenarios.
@@ IV-A Social Navigation Scenario
Consider a robot navigating next to $n\geq 1$ human agents in a workspace $\mathcal{W}\subseteq\mathbb{R}^{2}$ with a set of static obstacles $\mathcal{W}_{obs}\subseteq\mathcal{W}$. The robot starts from an initial configuration $s_{R}$ and moves towards a goal $g_{R}$ by following a policy $\pi_{R}$ whereas humans are navigating from their initial configurations $s_{i}$ towards their goals, $g_{i}$ by following a policy $\pi_{i}$, $i\in\mathcal{N}$; agents’ goals are unknown to one another. The robot occupies an area $A_{R}\in\mathcal{W}$, and each human occupies an area $A_{i}\in\mathcal{W}$. The objective of the robot is to reach its destination while avoiding collisions and abiding by social norms. We define a social navigation scenario as a tuple:
$\mathcal{S}=\left(n,A_{R},A_{1:n},s_{R},g_{R},s_{1:n},g_{1:n},\pi_{1:n},\mathcal{W}_{obs}\right)\mbox{.}$ (1)
@@ IV-B Experiment Design
Scenario configurations. We first define a base scenario $\mathcal{S}_{b}$ with $n=15$, $A_{R},A_{i:n}=\pi(0.3)^{2}$, $s_{R}=(5,1)$, $g_{R}=(5,9)$, $\pi_{i:n}=$ ORCA, SFM, $v_{pref}=1.0$ m/s, $\mathcal{W}_{obs}=\{[0,w]\times[0,l]\}^{\mathsf{c}}=\{[0,10]\times[0,10]\}^{\mathsf{c}}$; $s_{1:n},g_{1:n}$ are sampled from passing and crossing. We then modify a single variable in each experiment, from low to high intensity:
@@
  • Policy Mixture: {SFM only, ORCA only, Mix 1, Mix 2, Mix 3}, where Mix 1 contains 8 ORCA and 7 SFM agents, Mix 2 contains 5 ORCA, 5 SFM, 2 CV, 3 Static agents, and Mix 3 contains 4 ORCA, 4 SFM, 4 CV, 3 Static agents, respectively. We add increasing numbers of inattentive agents to model Complex real-world scenarios in which the robot must navigate among both cooperative and uncooperative agents simultaneously [41].
  • Environment Width: {4.5, 4.0, 3.5, 3.0, 2.5, 2.0, 1.5}. which reflects the reported Width of most hallways (e.g., [39, 60, 17]) in Table. I.
@@
(b) Density
(e) Width
(h) Directionality
(k) Policy Mixture
Figure 2: Left to right: Scenarios of increasing Complexity for each of the Complexity factors considered (b-k). In the leftmost Directionality figure green agents are crossing and blue agents are passing, while different colors represent different policies in Policy Mixture.
@@
  • Social GAN Model Predictive Control (MPC-SGAN): An approach integrating a recent crowd motion prediction model [67] into an MPC framework, using the approach and implementation of Poddar et al. [16].
@@
Metrics. Based on literature  [10, 1, 2], we use the following metrics for performance, collected from Successful trials:
@@
Hypotheses. We expect that scenarios of higher Complexity will pose greater navigation challenges, and this will be reflected in significant performance drops across all algorithms. Additionally, as Density is often used as a proxy for Complexity as a whole, we anticipate that it will have the strongest correlation with observed performance drops. Based on these expectations, We formulate the following hypotheses:
  • H1. Increasing the intensity of each of the four Complexity factors (Density, Directionality, Environment Geometry, Policy Mixture) independently decreases performance with respect to (w.r.t.) collected metrics.
@@ IV-C Implementation Details
We generate 500 scenarios for each condition within each experiment and report metrics as averages across all 500 trials. We fix the random seed to ensure each method experiences the same scenarios. To simulate continuous crowd motion, agents each have a precomputed sequence of goals sampled according to the Directionality of the scenario, and begin moving to their next goal upon reaching their current one. For MPC-SGAN and MPPI-SGAN, we used a checkpoint pretrained on the Zara portion of the UCY dataset [68]. RGL was retrained in 5, 10, and 15 agent scenarios with ORCA, SFM, and ORCA/SFM agent policies, however the training only resulted in a collision-avoiding policy in the 5 ORCA agent setting. For the ORCA and SFM methods, we used the CrowdNav default parameter settings. The reactive planner uses the same action space of Chen et al. [64], and selects the collision free action that minimizes its distance to the goal in the following timestep. All methods handle static obstacle collisions by setting the component of the action in the direction of the colliding obstacle to 00. For more specific details of the training procedure for RGL, and parameter settings and tuning for the MPC and MPPI methods, as well as our enhanced version of CrowdNav, see our GitHub site at https://github.com/fluentrobotics/ComplexityNav. Videos of our experiments can be found at https://youtu.be/-ir12VoSCkY.